# Exploring Pre-trained Text-to-Video Diffusion Models for Referring Video Object Segmentation

### v1

@@ Abstract
In this paper, we explore the visual representations produced from a pre-trained text-to-video (T2V) diffusion model for video understanding tasks. We hypothesize that the latent representation learned from a pretrained generative T2V model encapsulates rich semantics and coherent temporal correspondences, thereby naturally facilitating video understanding. Our hypothesis is validated through the classic referring video object segmentation (R-VOS) task. We introduce a novel framework, termed “VD-IT”, tailored with dedicatedly designed components built upon a fixed pretrained T2V model. Specifically, VD-IT uses textual information as a conditional input, ensuring semantic consistency across time for precise temporal instance matching. It further incorporates image tokens as supplementary textual inputs, enriching the feature set to generate detailed and nuanced masks.Besides, instead of using the standard Gaussian noise, we propose to predict the video-specific noise with an extra noise prediction module, which can help preserve the feature fidelity and elevates segmentation quality. Through extensive experiments, we surprisingly observe that fixed generative T2V diffusion models, unlike commonly used video backbones (e.g., Video Swin Transformer) pretrained with discriminative image/video pre-tasks, exhibit better potential to maintain semantic alignment and temporal consistency. On existing standard benchmarks, our VD-IT achieves highly competitive results, surpassing many existing state-of-the-art methods. The code will be available at https://github.com/buxiangzhiren/VD-IT
@@ 1 Introduction
However, the potential of pre-trained text-to-video diffusion models [2, 30, 58, 42] in video understanding tasks remains much less explored. Unlike image understanding, which primarily deals with single images, video comprehension necessitates the consideration of both spatial information within each frame and temporal information across all video frames, making it more challenging. Adhering to the principle of “what I cannot create, I do not understand”, we hypothesize that a pre-trained text-to-video diffusion model capable of generating coherent, high-quality video sequences based on text prompts inherently possesses sufficient requisite knowledge to help video understanding tasks.
@@ 4.1 Video Feature Extraction
More specifically, to extract visual tokens from video frames, we extract vectors using the CLIP [33] vision model to get embedding vectors $\{p_{v,t}\}_{t=1}^{T},p_{v,t}\in\mathbb{R}^{N_{p}\times C}$, where $N_{p}$ denotes the number of tokens and $C$ denotes the channel dimension. For referring text tokens, we utilize the text encoder provided by the text-to-video diffusion model to get the vectors $p_{e}\in\mathbb{R}^{L\times C}$. We use text token vectors guided image token projection as final prompt, which is formulate by
@@ 5.1 Comparison with State-of-the-Art Methods
$\mathcal{J}\&\mathcal{F}$
$\mathcal{J}$
$\mathcal{F}$
FPS
$\mathcal{J}\&\mathcal{F}$
$\mathcal{J}$
$\mathcal{F}$
CMSA [53] ResNet-50
36.4
34.8
38.1
40.2
36.9
43.5
URVOS [38] ResNet-50
47.2
45.3
49.2
51.5
47.3
56.0
CMPC-V [25] I3D
47.5
45.6
49.3
PMINet [9] ResNeSt-101
53.0
51.5
54.5
YOFO [17] ResNet-50
48.6
47.5
49.7
10
53.3
48.8
57.8
LBDT [8] ResNet-50
49.4
48.2
50.6
54.3
MLRL [46] ResNet-50
49.7
48.4
51.0
52.8
50.0
55.4
MTTR [3] Video-Swin-T
55.3
54.0
56.6
MANet [5] Video-Swin-T
55.6
54.8
56.5
ReferFormer [48] Video-Swin-T
56.0
54.8
57.3
50
SgMg [29] Video-Swin-T
58.9
57.7
60.0
65
56.7
53.3
60.0
SgMg* [29] Video-Swin-B
61.6
59.7
63.5
40
VD-IT (Ours) Video Diffusion
$\mathbf{64.8}$
$\mathbf{63.1}$
$\mathbf{66.6}$
21
$\mathbf{63.0}$
$\mathbf{59.9}$
$\mathbf{66.1}$
@@
ReferFormer [48] Video-Swin-T
59.4
58.0
60.9
50
59.6
56.5
62.7
SgMg [29] Video-Swin-T
62.0
60.4
63.5
65
61.9
59.0
64.8
ReferFormer [48] Video-Swin-B
62.9
61.3
64.6
33
61.1
58.1
64.1
OnlineRefer [47] Video-Swin-B
62.9
61.0
64.7
62.4
59.1
65.6
SgMg [29] Video-Swin-B
65.7
63.9
67.4
40
63.3
60.6
66.0
VD-IT (Ours) Video Diffusion
$\mathbf{67.2}$
$\mathbf{65.3}$
$\mathbf{69.1}$
21
$\mathbf{69.4}$
$\mathbf{66.2}$
$\mathbf{72.6}$
@@
mAP
Overall
Mean
mAP
Overall
Mean
Hu et al. [14] VGG-16
13.2
47.4
35.0
17.8
54.6
52.8
Gavrilyuk et al. [12] I3D
19.8
53.6
42.1
23.3
54.1
54.2
ACAN [41] I3D
27.4
60.1
49.0
28.9
57.6
58.4
CMPC-V [25] I3D
40.4
65.3
57.3
34.2
61.6
61.7
ClawCraneNet [22] ResNet-50/101
63.1
59.9
64.4
65.6
MTTR [3] Video-Swin-T
46.1
72.0
64.0
39.2
70.1
69.8
ReferFormer [48] Video-Swin-T
52.8
77.6
69.6
42.2
71.9
71.0
SgMg [29] Video-Swin-T
56.1
78.0
70.4
44.4
72.8
71.7
ReferFormer [48] Video-Swin-B
55.0
78.6
70.3
43.7
73.0
71.8
SgMg [29] Video-Swin-B
58.5
79.9
72.0
45.0
73.7
72.5
VD-IT (Ours) Video Diffusion
$\mathbf{61.4}$
$\mathbf{81.5}$
$\mathbf{73.2}$
$\mathbf{46.5}$
$\mathbf{74.4}$
$\mathbf{73.4}$
@@ 5.2 Ablation Study
Components Performance
Image-Cond
Text-Cond
NP
$\mathcal{J}\&\mathcal{F}$
$\mathcal{J}$
$\mathcal{F}$
$\checkmark$
59.7
57.9
61.6
$\checkmark$
61.9
60.1
63.7
$\checkmark$
$\checkmark$
63.8
62.0
65.5
$\checkmark$
$\checkmark$
$\checkmark$
$\mathbf{64.8}$
$\mathbf{63.1}$
$\mathbf{66.6}$
Method
Ref.
Ref.+
Ref.g
MaIL [21]
70.1
62.2
62.5
CRIS [45]
70.5
62.3
59.9
RefTR [18]
70.6
LAVT [52]
72.7
62.1
61.2
VLT [7]
73.0
63.5
63.5
SgMg [29]
76.3
66.4
70.0
VD-IT (Ours)
$\mathbf{76.7}$
$\mathbf{66.5}$
$\mathbf{70.3}$
@@
Method IoU Diff.
Inter. 1
Inter. 5
SgMg [29]
7.24
11.15
VD-I
6.52
9.43
VD-T
5.41
8.25
VD-IT
$\mathbf{5.19}$
$\mathbf{7.89}$
@@
Image-Cond
Text-Cond
NP
$\checkmark$
47.26%
$\checkmark$
44.54%
$\checkmark$
$\checkmark$
46.71%
$\checkmark$
$\checkmark$
$\checkmark$
$\mathbf{47.36\%}$
@@
[Uncaptioned image] Figure 4: Temporal Semantic Consistency. Averaged over 1,000 samples from Ref-Youtube-VOS, the cosine similarity between the Region of Interest (RoI) features of the initial frame and the following seven frames is reported.
[Uncaptioned image] Figure 5: Robustness against light noise. We modify the brightness of various frames randomly and compare the IoU of segmentation results under changing lighting conditions. The results are reported on Ref-Youtube-VOS.
@@
Visualization Results. In Figure 6, we further provide some visual comparison results with the state-of-the-art method SgMg [29]. These include scenarios with significant variations in motion and pose, drastic changes in lighting, and instances where the object is partially obscured. From the visual results, we can clearly see that our proposed VD-IT can segment and track the referred object in a more temporally consistent way than the baseline method.
@@ Appendix 0.B Additional Visualization Results
Method
Fusion
$\mathcal{J}\&\mathcal{F}$
$\mathcal{J}$
$\mathcal{F}$
VD-I
59.7
57.9
61.6
VD-T
61.9
60.1
63.7
VD-IT
Concat.
62.4
60.8
64.0
VD-IT
Attn.
$\mathbf{64.8}$
$\mathbf{63.1}$
$\mathbf{66.6}$
Step
$\mathcal{J}\&\mathcal{F}$
$\mathcal{J}$
$\mathcal{F}$
1
$\mathbf{64.8}$
$\mathbf{63.1}$
$\mathbf{66.6}$
5
63.4
61.6
65.1
10
63.5
61.7
65.1
50
63.1
61.3
64.9
100
62.7
60.8
64.5

### v2

@@ Abstract
In this paper, we explore the visual representations produced from a pre-trained text-to-video (T2V) diffusion model for video understanding tasks. We hypothesize that the latent representation learned from a pretrained generative T2V model encapsulates rich semantics and coherent temporal correspondences, thereby naturally facilitating video understanding. Our hypothesis is validated through the classic referring video object segmentation (R-VOS) task. We introduce a novel framework, termed “VD-IT”, tailored with dedicatedly designed components built upon a fixed pretrained T2V model. Specifically, VD-IT uses textual information as a conditional input, ensuring semantic consistency across time for precise temporal instance matching. It further incorporates image tokens as supplementary textual inputs, enriching the feature set to generate detailed and nuanced masks. Besides, instead of using the standard Gaussian noise, we propose to predict the video-specific noise with an extra noise prediction module, which can help preserve the feature fidelity and elevates segmentation quality. Through extensive experiments, we surprisingly observe that fixed generative T2V diffusion models, unlike commonly used video backbones (e.g., Video Swin Transformer) pretrained with discriminative image/video pre-tasks, exhibit better potential to maintain semantic alignment and temporal consistency. On existing standard benchmarks, our VD-IT achieves highly competitive results, surpassing many existing state-of-the-art methods. The code is available at https://github.com/buxiangzhiren/VD-IT.
@@ 1 Introduction
However, the potential of pre-trained text-to-video diffusion models [2, 30, 62, 42] in video understanding tasks [46, 45, 44, 43] remains much less explored. Unlike image understanding, which primarily deals with single images, video comprehension necessitates the consideration of both spatial information within each frame and temporal information across all video frames, making it more challenging. Adhering to the principle of “what I cannot create, I do not understand”, we hypothesize that a pre-trained text-to-video diffusion model capable of generating coherent, high-quality video sequences based on text prompts inherently possesses sufficient requisite knowledge to help video understanding tasks.
@@ 4.1 Video Feature Extraction
More specifically, to extract visual tokens from video frames, we extract vectors using the CLIP [33] vision model to get embedding vectors $\{p_{v,t}\}_{t=1}^{T},p_{v,t}\in\mathbb{R}^{N_{p}\times C}$, where $N_{p}$ denotes the number of tokens and $C$ denotes the channel dimension. For referring text tokens, we utilize the text encoder provided by the text-to-video diffusion model to get the vectors $p_{e}\in\mathbb{R}^{L\times C}$. We use text token vectors guided image token projection as final prompt, which is formulated by
@@ 5.1 Comparison with State-of-the-Art Methods
[…]
CMSA [57] ResNet-50 36.4 34.8 38.1 - 40.2 36.9 43.5
URVOS [38] ResNet-50 47.2 45.3 49.2 - 51.5 47.3 56.0
CMPC-V [25] I3D 47.5 45.6 49.3 - - - -
PMINet [9] ResNeSt-101 53.0 51.5 54.5 - - - -
YOFO [17] ResNet-50 48.6 47.5 49.7 10 53.3 48.8 57.8
LBDT [8] ResNet-50 49.4 48.2 50.6 - 54.3 - -
MLRL [50] ResNet-50 49.7 48.4 51.0 - 52.8 50.0 55.4
MTTR [3] Video-Swin-T 55.3 54.0 56.6 - - - -
MANet [5] Video-Swin-T 55.6 54.8 56.5 - - - -
ReferFormer [52] Video-Swin-T 56.0 54.8 57.3 50 - - -
SgMg [29] Video-Swin-T 58.9 57.7 60.0 65 56.7 53.3 60.0
SgMg* [29] Video-Swin-B 61.6 59.7 63.5 40 - - -
VD-IT (Ours) Video Diffusion $\mathbf{64.8}$ $\mathbf{63.1}$ $\mathbf{66.6}$ 21 $\mathbf{63.0}$ $\mathbf{59.9}$ $\mathbf{66.1}$
@@
ReferFormer [52] Video-Swin-T 59.4 58.0 60.9 50 59.6 56.5 62.7
SgMg [29] Video-Swin-T 62.0 60.4 63.5 65 61.9 59.0 64.8
ReferFormer [52] Video-Swin-B 62.9 61.3 64.6 33 61.1 58.1 64.1
OnlineRefer [51] Video-Swin-B 62.9 61.0 64.7 - 62.4 59.1 65.6
SgMg [29] Video-Swin-B 65.7 63.9 67.4 40 63.3 60.6 66.0
VD-IT (Ours) Video Diffusion $\mathbf{66.5}$ $\mathbf{64.4}$ $\mathbf{68.5}$ 21 $\mathbf{69.4}$ $\mathbf{66.2}$ $\mathbf{72.6}$
@@
mAP Overall Mean mAP Overall Mean
Hu et al. [14] VGG-16 13.2 47.4 35.0 17.8 54.6 52.8
Gavrilyuk et al. [12] I3D 19.8 53.6 42.1 23.3 54.1 54.2
ACAN [41] I3D 27.4 60.1 49.0 28.9 57.6 58.4
CMPC-V [25] I3D 40.4 65.3 57.3 34.2 61.6 61.7
ClawCraneNet [22] ResNet-50/101 - 63.1 59.9 - 64.4 65.6
MTTR [3] Video-Swin-T 46.1 72.0 64.0 39.2 70.1 69.8
ReferFormer [52] Video-Swin-T 52.8 77.6 69.6 42.2 71.9 71.0
SgMg [29] Video-Swin-T 56.1 78.0 70.4 44.4 72.8 71.7
ReferFormer [52] Video-Swin-B 55.0 78.6 70.3 43.7 73.0 71.8
SgMg [29] Video-Swin-B 58.5 79.9 72.0 45.0 73.7 72.5
VD-IT (Ours) Video Diffusion $\mathbf{61.4}$ $\mathbf{81.5}$ $\mathbf{73.2}$ $\mathbf{46.5}$ $\mathbf{74.4}$ $\mathbf{73.4}$
@@ 5.2 Ablation Study
Components Performance
[…]
$\checkmark$ 59.7 57.9 61.6
$\checkmark$ 61.9 60.1 63.7
$\checkmark$ $\checkmark$ 63.8 62.0 65.5
$\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{64.8}$ $\mathbf{63.1}$ $\mathbf{66.6}$
@@
Method Ref. Ref.+ Ref.g
MaIL [21] 70.1 62.2 62.5
CRIS [49] 70.5 62.3 59.9
RefTR [18] 70.6 - -
LAVT [56] 72.7 62.1 61.2
VLT [7] 73.0 63.5 63.5
SgMg [29] 76.3 66.4 70.0
VD-IT (Ours) $\mathbf{76.7}$ $\mathbf{66.5}$ $\mathbf{70.3}$
@@
Method IoU Diff.
Inter. 1 Inter. 5
SgMg [29] 7.24 11.15
VD-I 6.52 9.43
VD-T 5.41 8.25
VD-IT $\mathbf{5.19}$ $\mathbf{7.89}$
@@
Image-Cond Text-Cond NP
$\checkmark$ 47.26%
$\checkmark$ 44.54%
$\checkmark$ $\checkmark$ 46.71%
$\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{47.36\%}$
@@
Figure 4: Temporal Semantic Consistency. Averaged over 1,000 samples from Ref-Youtube-VOS, the cosine similarity between the Region of Interest (RoI) features of the initial frame and the following seven frames is reported.
Figure 5: Robustness against light noise. We modify the brightness of various frames randomly and compare the IoU of segmentation results under changing lighting conditions. The results are reported on Ref-Youtube-VOS.
@@
Visualization Results. In Figure 6, we further provide some visual comparison results with the state-of-the-art method SgMg [29]. From the visual results, we can clearly see that our proposed VD-IT can segment and track the referred object in a more temporally consistent way than the baseline method.
@@ Appendix 0.B Additional Visualization Results
[…]
1 $\mathbf{64.8}$ $\mathbf{63.1}$ $\mathbf{66.6}$
5 63.4 61.6 65.1
10 63.5 61.7 65.1
50 63.1 61.3 64.9
100 62.7 60.8 64.5